How do u solve x = 2 and how is it show on a graph.
Answer by AnlytcPhil(1116)

Hw do u solve x = 2
and hw is it show on a graph..........?????? please help
Consider the equation to have a 0y term, like this
x + 0y = 2
Now substitute a few arbitrary numbers for y, say,
y=1, y=-3, y=4, and y=-2. Substituting y=1
x + 0(1) = 2
x = 2, so now you have the point (2,1)
Substituting y=-3
x + 0(-3) = 2
x = 2, so now you have the point (2,-3)
Substituting y=4
x + 0(4) = 2
x = 2, so now you have the point (2,4)
Substituting y=-2
x + 0(-2) = 2
x = 2, so now you have the point (2,-2)
Plot those points, connect them and you'll get a vertical line
parallel to the y-axis and two units right of it.
Edwin